# Modeling Microenvironment Trajectories on Spatial Transcriptomics with NicheFlow

### v1

@@ D.9 Wasserstein metrics
To complement the local structural assessments with GW and FGW, we additionally include global evaluation metrics that are widely adopted in spatial transcriptomics [48, 49, 50]. Specifically, we report the 1-Wasserstein ($\mathcal{W}_{1}$) and 2-Wasserstein ($\mathcal{W}_{2}$) distances between predicted and real cell distributions, computed separately over spatial coordinates and gene expression features. These metrics are calculated at the level of individual cell types and averaged across all timepoints and generated samples. This provides a more global perspective on how well the model reconstructs tissue-wide spatial and transcriptional distributions.
@@ F.2 Dataset and microenvironments preprocessing
All datasets used in our study underwent a preprocessing procedure appropriate for spatial transcriptomic analysis, involving total count normalization, logarithmic transformation, and principal component analysis (PCA). For the mouse embryonic development [20] and axolotl brain development [42] datasets, total count normalization, and logarithmic transformation had already been applied; we, therefore, performed PCA ourselves on the transformed data. For the mouse brain aging dataset [43], we applied all three steps: we first normalized raw count matrices so that each cell had the same total expression, followed by a natural logarithm transformation of the form $\log(X+1)$ to stabilize variance and mitigate the influence of large values. We then computed PCA on the log-transformed data.

### v2

@@ D.9 Wasserstein metrics
To complement the local structural assessments with GW and FGW, we additionally include global evaluation metrics that are widely adopted in spatial transcriptomics [wasserstein-ex-trajectorynet, wasserstein-ex-mioflow, wasserstein-ex-deepruot]. Specifically, we report the 1-Wasserstein ($\mathcal{W}_{1}$) and 2-Wasserstein ($\mathcal{W}_{2}$) distances between predicted and real cell distributions, computed separately over spatial coordinates and gene expression features. These metrics are calculated at the level of individual cell types and averaged across all timepoints and generated samples. This provides a more global perspective on how well the model reconstructs tissue-wide spatial and transcriptional distributions.
@@ F.2 Dataset and microenvironments preprocessing
All datasets used in our study underwent a preprocessing procedure appropriate for spatial transcriptomic analysis, involving total count normalization, logarithmic transformation, and principal component analysis (PCA). For the mouse embryonic development [klein2025mapping] and axolotl brain development [axolotl] datasets, total count normalization, and logarithmic transformation had already been applied; we, therefore, performed PCA ourselves on the transformed data. For the mouse brain aging dataset [Sun2025aging], we applied all three steps: we first normalized raw count matrices so that each cell had the same total expression, followed by a natural logarithm transformation of the form $\log(X+1)$ to stabilize variance and mitigate the influence of large values. We then computed PCA on the log-transformed data.